# Detecting Masquerade Attacks in Controller Area Networks Using Graph Machine Learning

### v1

@@ V-C Metrics Summary
We compare the effectiveness of two detection settings: using embeddings only and embeddings combined with time series features. The evaluation focuses on the AUC-ROC metric (see Section IV F) to compare the performance of these settings against various attack scenarios.
@@
Table II summarizes results across heatmaps. We place attack categories in the columns and detection settings in the rows. This means that each cell in this table shows the summary statistics from the heatmaps for a specific detection setting and attack. We focus on summary statistics including mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), and maximum ($\max$). In addition, we also compute average ($\overline{\max}$) and standard deviation ($\sigma_{\max}$) of the maximum values for each of the attack detection settings and attack types in the ROAD dataset. They are displayed as last rows and columns in the table.
@@
Graph embeddings combined with time series features generally outperform the graph embeddings-only setting across all attack types. Specifically, for ‘correlated signal attack’, the graph embeddings with time series features setting achieves a mean AUC-ROC ($\mu$) of 0.96 with a standard deviation ($\sigma$) of 0.02, compared to the embeddings-only setting with a mean AUC-ROC of 0.94 and a standard deviation of 0.02. For ‘max engine coolant temperature attack’, the embeddings with time series setting has a mean AUC-ROC of 0.94 and a standard deviation of 0.04, which is higher than the graph embeddings-only setting with a mean AUC-ROC of 0.91 and a standard deviation of 0.04. In the case of ‘max speedometer attack’, the mean AUC-ROC for the graph embeddings with time series setting is 0.94 with a standard deviation of 0.03. In contrast, the embeddings-only setting achieves a mean AUC-ROC of 0.91 with a standard deviation of 0.03. For ‘reverse light off attack’, the graph embeddings with time series setting achieves a mean AUC-ROC of 0.96 with a standard deviation of 0.02, while the graph embeddings-only setting has a mean AUC-ROC of 0.93 and a standard deviation of 0.02. Finally, for ‘reverse light on attack’, the mean AUC-ROC for the graph embeddings with time series setting is 0.94 with a standard deviation of 0.03, compared to the graph embeddings-only setting with a mean AUC-ROC of 0.92 and a standard deviation of 0.03.
In ‘correlated signal attack’, the graph embeddings with time series setting achieves a median ($\eta$) of 0.96, a minimum ($\min$) of 0.92, and a maximum ($\max$) of 0.98. These metrics provide further insights into the detection performance by highlighting the central tendency and range of the results. The average ($\overline{\max}$) and standard deviation ($\sigma_{\max}$) of the maximum values obtained by each attack and experiment type are also reported. These values illustrate the peak performance and consistency of the settings, respectively. For instance, for ‘max engine coolant temperature attack’, the average maximum AUC-ROC ($\overline{\max}$) is 0.98 for the embeddings only setting and 0.99 for the embeddings with time series setting. The standard deviation ($\sigma_{\max}$) for both settings is 0.00, demonstrating consistent peak performance across different experiments.
To further validate the differences in detection effectiveness of both settings, we employed the Mann-Whitney $\mathrm{U}$ test and the Kolmogorov-Smirnov $(\mathrm{KS}$) test. The Mann-Whitney $\mathrm{U}$ test [75] is a non-parametric statistical test that evaluates whether there is a difference between two independent samples. The $\mathrm{KS}$ test [76] is another non-parametric test that assesses whether two samples come from the same distribution. In our analysis, we compared the AUC-ROC value distributions of the two settings: one using only embeddings and the other combining these with time series features. Here, the null hypothesis is that the AUC-ROC values for the setting combining embeddings with time series features are less than or equal to those for the setting using only embeddings, while the alternative hypothesis is that the AUC-ROC values for the setting combining embeddings with time series features are greater than those for the setting using only embeddings. A low $p$-value indicates a significant difference between these two settings. We focused on a significance level of 0.05. Table III shows significant differences between the observed distributions for all masquerade attacks. In each of the attacks, the Mann-Whitney U and $\mathrm{KS}$ tests produce low $p$-values, indicating a significant deviation from the expected distribution. This supports our hypothesis of relying on annotated MSGs to effectively detect masquerade attacks at various windows and offset combinations. Consequently, we reject the null hypothesis for all tested masquerade attacks, confirming that the combined setting performs significantly better.
TABLE III: Mann-Whitney U and $\mathrm{KS}$ tests results for different masquerade attacks.
Attack name Mann-Whitney Kolmogorov- Smirnov
U Statistic P-value Statistic P-value
Correlated signal 11337.0 5.55e-16 0.462 3.81e-12
Max speedometer 10082.0 7.98e-09 0.328 2.34e-06
Reverse light off 11033.0 4.98e-14 0.429 1.73e-10
Reverse light on 10818.0 9.84e-13 0.429 1.73e-10
Max engine coolant 9864.0 8.01e-08 0.328 2.34e-06
Figure 9: AUC-ROC across different window size and offset combinations for masquerade attacks detected by the baseline model using only graph embeddings.
Figure 10: AUC-ROC across different window size and offset combinations for masquerade attacks detected by incorporating both graph embeddings and time series features.
V-D Evaluation Based on Sample Windows
We investigate the effect of sample-based windows in our detection framework’s performance, as opposed to time-based windows. Our aim is to understand the impact of varying the number of CAN message samples in each sliding window on masquerade attack detection. In this context, time-based windows refer to windows defined by a fixed duration, whereas sample-based windows are defined by a fixed number of CAN messages. We experiment with different window sizes and offsets to find the optimal configurations for detecting anomalies.
Figs. 10 and 10 are heatmaps that provide visual comparisons of the AUC-ROC values for different $\omega$ and $\delta$ combinations for both settings. These heatmaps show that the combined setting generally outperforms the embeddings-only setting, especially at certain window sizes and offsets. However, the optimal configuration varies significantly depending on the attack type, emphasizing the importance of careful parameter selection in practical applications. For instance, in the ‘correlated signal attack’, the combined setting achieves a higher AUC-ROC value of 0.84 with a $\omega$ of 100 and an $\delta$ of 100, compared to the embeddings-only setting, which peaks at 0.79 with the same configuration. Similarly, in the ‘max engine coolant temperature attack’, the combined setting reaches an AUC-ROC value of 0.97 with a $\omega$ of 350 and an $\delta$ of 300, outperforming the embeddings-only setting, which attains a maximum of 0.89 with a $\omega$ of 200 and an $\delta$ of 200. In the ‘reverse light off attack’, the combined setting shows superior performance with an AUC-ROC value of 0.81 with a $\omega$ of 400 and an $\delta$ of 400, while the embeddings-only setting reaches 0.76 with the same configuration. Note that in the remaining attacks (i.e., ‘max speedometer attack’ and ‘reverse light on attack’), combining graph embeddings with time series features allows the classifier to obtain higher AUC-ROC values but requires a lower number of samples.
Table V presents the summary statistics for all masquerade attacks in the ROAD dataset. For both settings, the highest and lowest performances occur in the ‘max engine coolant temperature attack’ scenario, with varying window sizes and offsets. In the graph embeddings-only setting, we achieve a mean AUC-ROC ($\mu$) of 0.61 with a standard deviation ($\sigma$) of 0.06. When we combine graph embeddings with time series features, the mean AUC-ROC ($\mu$) improves to 0.68, with a standard deviation ($\sigma$) of 0.05. Our findings back previous results that the integration of time series features with graph embeddings generally enhances the model’s performance.
The granular window approach proves effective with performance varying significantly across different attack types and window configurations. Individual attack scenarios reveal interesting patterns. For instance, ‘correlated signal attack’ shows consistent improvement when time series features are added, with the mean AUC-ROC increasing from 0.68 to 0.74. The ‘max engine coolant temperature attack’ exhibits the highest variability in performance, indicating high sensitivity to $\omega$ and $\delta$ selection. The ‘max speedometer attack’ and ‘reverse light on attack’ prove more challenging to detect, with lower mean AUC-ROC values in both settings. The ‘reverse light off attack’ shows moderate improvement with the addition of time series features. We observed that the average $\omega$ increased slightly from 240 to 260 samples when incorporating time series features, while the average $\delta$ remained constant at 220 samples. Despite this minor increase in $\omega$, the combined approach consistently demonstrated superior detection performance across various attack scenarios, as evidenced by the improved AUC-ROC values. For instance, in ‘correlated signal attack’, the AUC-ROC improved from 0.79 to 0.84, while in ‘max engine coolant temperature attack’, it increased from 0.89 to 0.97. These results suggest that the integration of time series features with graph embeddings enhances the model’s ability to detect anomalies, even when requiring slightly larger sample windows. This improvement in detection capability outweighs the small increase in computational requirements, further validating the effectiveness of our combined approach in various window configurations.
We also found the distributions of AUC-ROC values to be significantly different between the two settings. The Mann-Whitney $\mathrm{U}$ and $\mathrm{KS}$ tests confirmed this difference, indicating the superior detection performance of the combined setting. Table IV summarizes the results, showing significant differences across all tested masquerade attacks. For instance, ‘correlated signal attack’ yielded a Mann-Whitney U statistic of 1069.5 with a $p$-value of 9.84e-07, and a KS statistic of 0.58 with a $p$-value of 2.61e-06. Similar trends were observed for other attacks, such as ‘max engine coolant’ and ‘max speedometer’ attacks, confirming the effectiveness of incorporating time series features with graph embeddings. The results for each attack type consistently indicate that the combined setting outperforms the embeddings-only setting, validating the improvements seen in the AUC-ROC metrics.
TABLE IV: Mann-Whitney U and Kolmogorov-Smirnov test results for different masquerade attacks in the sample-based window experiment.
Attack name Mann-Whitney Kolmogorov- Smirnov
U Statistic P-value Statistic P-value
Correlated signal 1069.5 9.84e-07 0.58 2.61e-06
Max engine coolant 811.5 0.033 0.33 0.018
Max speedometer 1085.0 4.20e-07 0.56 9.30e-06
Reverse light off 1040.5 4.84e-06 0.50 9.36e-05
Reverse light on 1112.0 8.35e-08 0.61 6.75e-07
TABLE V: Summary statistics for the evaluation metrics based on sample-based windows. We show the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), and maximum ($\max$) of the AUC ROC for each setting across different attack scenarios, as well as the average ($\overline{\max}$) and standard deviation ($\sigma_{\max}$) of the maximum values obtained by each attack and experiment type, either embeddings-only or embeddings with time series information.
@@ V-D Evaluation Based on Sample Windows
Experiment Correlated signal attack Max engine coolant temperature attack Max speedometer attack Reverse light off attack Reverse light on attack Experiment max
Graph embeddings- only $\mu=0.68$ $\sigma=0.05$ $\eta=0.68$ min = 0.54 (400, 400) max = 0.79 (100, 100) $\mu=0.63$ $\sigma=0.19$ $\eta=0.69$ min = 0.17 (400, 300) max = 0.89 (200, 200) $\mu=0.56$ $\sigma=0.05$ $\eta=0.56$ min = 0.45 (300, 250) max = 0.65 (350, 250) $\mu=0.63$ $\sigma=0.05$ $\eta=0.64$ min = 0.52 (300, 200) max = 0.76 (400, 400) $\mu=0.55$ $\sigma=0.05$ $\eta=0.56$ min = 0.45 (350, 350) max = 0.68 (150, 150) $\overline{\text{max}}=0.75$ $\sigma_{\text{max}}=0.10$
Graph embeddings + Time series features $\mu=0.74$ $\sigma=0.05$ $\eta=0.75$ min = 0.60 (400, 400) max = 0.84 (100, 100) $\mu=0.70$ $\sigma=0.19$ $\eta=0.76$ min = 0.24 (200, 150) max = 0.97 (350, 300) $\mu=0.63$ $\sigma=0.04$ $\eta=0.64$ min = 0.52 (300, 250) max = 0.70 (300, 150) $\mu=0.69$ $\sigma=0.05$ $\eta=0.69$ min = 0.58 (50, 50) max = 0.81 (400, 400) $\mu=0.62$ $\sigma=0.04$ $\eta=0.61$ min = 0.52 (350, 350) max = 0.72 (150, 150) $\overline{\text{max}}=0.81$ $\sigma_{\text{max}}=0.11$
Attack max $\overline{\text{max}}=0.81$ $\sigma_{\text{max}}=0.04$ $\overline{\text{max}}=0.93$ $\sigma_{\text{max}}=0.06$ $\overline{\text{max}}=0.68$ $\sigma_{\text{max}}=0.04$ $\overline{\text{max}}=0.79$ $\sigma_{\text{max}}=0.04$ $\overline{\text{max}}=0.70$ $\sigma_{\text{max}}=0.03$

### v2

@@ V-C Metrics Summary
We compare the effectiveness of the two previously defined configurations and include a time series-only variant as an ablation baseline to assess the individual contribution of temporal features.
@@
Table III summarizes results across heatmaps. Attack categories are listed in the columns, and detection settings in the rows. Each cell in this table reports summary statistics from the heatmaps for a specific setting and attack type. We focus on metrics including mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), and maximum ($\max$). Additionally, we compute the average ($\overline{\max}$) and standard deviation ($\sigma_{\max}$) of the maximum values obtained by each setting and attack type, which are shown in the last row and column of the table.
@@
Graph embeddings combined with time series features consistently outperform both the graph embeddings-only and the time series-only baselines across all attack types. For example, for the correlated_signal_attack, the hybrid setting achieves a mean AUC-ROC ($\mu$) of 0.96 with a standard deviation ($\sigma$) of 0.02, compared to 0.94 and 0.02 for the graph-only baseline, and 0.82 and 0.05 for the time-series-only baseline. Similarly, for max_engine_coolant_attack, the hybrid setting achieves a mean AUC-ROC of 0.94 and standard deviation of 0.04, versus 0.91 and 0.04 for graph-only, and 0.79 and 0.06 for time-series-only. A comparable improvement is seen across max_speedometer_attack (0.94 vs. 0.91 vs. 0.85), reverse_light_off_attack (0.96 vs. 0.93 vs. 0.81), and reverse_light_on_attack (0.94 vs. 0.92 vs. 0.84), with all standard deviations remaining within similar bounds.
How do spoofed frames differ in feature domains?
In feature space, spoofed frames differ from benign traffic primarily through anomalies in structural transitions and temporal regularities [moriano2024benchmarkingunsupervisedonlineids]. Our evaluation shows that the fused feature representation enables near-perfect separation between benign and malicious traffic. Specifically, our hybrid model achieves a mean AUC-ROC of 0.99 across all masquerade attacks, indicating that the combined structural and temporal deviations are sufficient for discrimination, even when content and timing are superficially plausible. Notably, difficult-to-detect attacks such as max_engine_coolant, previously noted for weak signal-level deviation, are still effectively detected by our approach. This supports the conclusion that our hybrid features amplify subtle inconsistencies—particularly in message transition structure and short-term temporal signal patterns—that standalone models may miss. Prior CAN intrusion studies also suggest that such composite feature sets yield superior discrimination [cho2016fingerprinting, tariq2020cantransfer].
We now summarize median and maximum statistics for each configuration. In correlated_signal_attack, the hybrid setting yields a median ($\eta$) of 0.96, a minimum ($\min$) of 0.92, and a maximum ($\max$) of 0.99. These statistics capture the central tendency and spread of detection performance across all configurations. Notably, the time-series-only baseline consistently underperforms in terms of both average and peak metrics, which further validates the necessity of combining temporal and structural features for detecting stealthy attacks. Moreover, we observe that for max_engine_coolant_attack, the average maximum AUC-ROC ($\overline{\max}$) reaches 0.99 for the hybrid model, 0.98 for the graph-only baseline, and only 0.89 for the time-series-only variant, with a standard deviation of ($\sigma_{\max}$) = 0.00 or 0.01—highlighting both the performance gap and stability offered by the hybrid method across experimental trials.
To provide clearer insights into the IDS’s classification performance, we computed the false positive rate (FPR) and false negative rate (FNR) at a selected operating threshold derived from the receiver operating characteristic (ROC) curve [FAWCETT2006861]. Specifically, we selected the threshold that maximizes Youden’s $J$ index ($J=\text{TPR}-\text{FPR}$), a standard method for identifying the optimal balance between true positive and false positive rates in binary classification tasks [yang2025unflows]. This ensures a meaningful trade-off between sensitivity and specificity, supporting practical deployment. Table IV reports the average FPR and FNR across all attack types and $(\omega_{t},\delta_{t})$ combinations, highlighting consistent improvements—especially in FNR—when time series features are integrated.
TABLE IV: MEAN AND STANDARD DEVIATION OF FPR AND FNR AT THE OPTIMAL ROC-DERIVED THRESHOLD (YOUDEN’S $J$ INDEX). “GRAPH EMBEDDINGS” REFERS TO THE BASELINE; “+ TIME SERIES FEATURES” DENOTES THE ENHANCED SETTING.
@@
Attack Name Version FPR FNR (mean $\pm$ std) (mean $\pm$ std) correlated_signal Graph embeddings 0.04 $\pm$ 0.01 0.09 $\pm$ 0.02 + time series features 0.02 $\pm$ 0.01 0.07 $\pm$ 0.02 max_speedometer Graph embeddings 0.01 $\pm$ 0.01 1.20 $\pm$ 0.22 + time series features 0.00 $\pm$ 0.01 0.80 $\pm$ 0.17 max_engine_coolant Graph embeddings 0.02 $\pm$ 0.01 0.27 $\pm$ 0.05 + time series features 0.01 $\pm$ 0.01 0.21 $\pm$ 0.05 reverse_light_off Graph embeddings 0.06 $\pm$ 0.01 0.01 $\pm$ 0.01 + time series features 0.04 $\pm$ 0.01 0.01 $\pm$ 0.01 reverse_light_on Graph embeddings 0.01 $\pm$ 0.01 0.87 $\pm$ 0.19 + time series features 0.00 $\pm$ 0.01 0.69 $\pm$ 0.14
Table V summarizes TTW results, including mean, median, standard deviation, minimum, and maximum processing times, with corresponding optimal $\delta_{t}$ and $\omega_{t}$ combinations shown in parentheses. Table V demonstrates that our hybrid approach achieves reasonable computational overhead across attack types. For instance, the correlated_signal_attack achieves a mean TTW of 2.36s, with a best-case performance as low as 0.024s, highlighting feasibility for real-time application with further optimization.
TABLE V: TTW SUMMARY ACROSS ATTACK TYPES WITH MIN/MAX CONFIGURATIONS.
Attack Name Mean (s) Std (s) Median (s) Min TTW (s) Max TTW (s) correlated_signal 2.36 1.21 2.44 0.024 (12.0, 1.0) 6.84 (10.0, 1.0) max_speedometer 2.67 1.29 2.65 0.005 (13.0, 4.0) 20.82 (13.0, 4.0) max_engine_coolant 2.30 1.16 2.30 0.237 (10.0, 5.0) 4.89 (15.0, 9.0) reverse_light_off 2.60 1.29 2.59 0.027 (10.0, 1.0) 6.22 (15.0, 14.0) reverse_light_on 2.80 1.36 2.93 0.005 (13.0, 1.0) 8.53 (9.0, 2.0)
Table VI presents the AUC-ROC and corresponding TTW for selected $(\omega_{t},\delta_{t})$ configurations across all attack types. Each row reflects a configuration that balances high detection accuracy with practical runtime cost. Note that for reverse_light_on, the second row highlights a lower-latency alternative with a slight reduction in AUC-ROC, illustrating the effect of window size and offset on real-time feasibility. We show that our hybrid approach consistently achieves near-perfect detection (0.99 AUC-ROC) across all masquerade attacks, while also revealing the varying computational overhead associated with different optimal configurations. For instance, the correlated_signal attack is detected with 0.99 AUC-ROC at 0.44s TTW, highlighting its efficiency. This comprehensive view of performance and runtime is critical for deploying IDS in resource-constrained vehicular environments.
TABLE VI: TRADE-OFF BETWEEN DETECTION ACCURACY AND RUNTIME EFFICIENCY FOR SELECTED $(\omega_{t},\delta_{t})$ CONFIGURATIONS.
Attack Name $\omega_{t}$ (s) $\delta_{t}$ (s) AUC-ROC TTW (s) correlated_signal 3 3 0.99 0.44 max_speedometer 8 8 0.99 1.85 max_engine_coolant 7 4 0.99 0.92 reverse_light_off 5 5 0.99 0.73 reverse_light_on 7 4 0.99 1.10 4 2 0.96 0.18
Figure 8 visualizes this trade-off by plotting AUC-ROC against TTW, annotated with the corresponding $(\omega_{t},\delta_{t})$ values. While most configurations achieve near-identical AUC-ROC scores (0.99), the associated TTW varies, reflecting the computational impact of wider temporal contexts. Note that the alternate reverse_light_on configuration, for example, reduces TTW from 1.10 s to 0.18 s with only a minor drop in AUC-ROC. We show the importance of selecting $(\omega_{t},\delta_{t})$ pairs that offer an optimal balance between detection accuracy and inference latency.
Figure 8: Trade-off between detection accuracy (AUC-ROC) and runtime cost (TTW) for selected $(\omega_{t},\delta_{t})$ configurations. Values in parentheses below each point indicate the $\omega_{t}$ and $\delta_{t}$ used in that configuration. The reverse_light_on (alt) point reflects an alternate low-latency configuration for the same attack, included to illustrate the practical trade-off.
To further validate the differences in detection effectiveness of both settings, we employed the Mann-Whitney $\mathrm{U}$ test and the Kolmogorov-Smirnov $(\mathrm{KS}$) test. The Mann-Whitney $\mathrm{U}$ test [mann1947test] is a non-parametric statistical test that evaluates whether there is a difference between two independent samples. The $\mathrm{KS}$ test [berger2014kolmogorov] is another non-parametric test that assesses whether two samples come from the same distribution. In our analysis, we compared the AUC-ROC value distributions of the two settings. Here, the null hypothesis is that the AUC-ROC values for the setting combining embeddings with time series features are less than or equal to those for the setting using only embeddings, while the alternative hypothesis is that the AUC-ROC values for the setting combining embeddings with time series features are greater than those for the setting using only embeddings. A low $p$-value indicates a significant difference between these two settings. We focused on a significance level of 0.05.
Table VII shows significant differences between the observed distributions for all masquerade attacks. In each of the attacks, the Mann-Whitney U and $\mathrm{KS}$ tests produce low $p$-values, indicating a significant deviation from the expected distribution. This supports our hypothesis of relying on annotated MSGs to effectively detect masquerade attacks at various windows and offset combinations. Consequently, we reject the null hypothesis for all tested masquerade attacks, confirming that the combined setting performs significantly better.
TABLE VII: MANN-WHITNEY U AND KOLMOGOROV-SMIRNOV TEST RESULTS FOR TIME-BASED WINDOWS.
Attack Name U Statistic $p$-value KS Statistic $p$-value correlated_signal 11337.0 5.55e-16 0.46 3.81e-12 max_speedometer 10082.0 7.98e-09 0.33 2.34e-06 max_engine_coolant 9864.0 8.01e-08 0.33 2.34e-06 reverse_light_off 11033.0 4.98e-14 0.43 1.73e-10 reverse_light_on 10818.0 9.84e-13 0.43 1.73e-10
V-D Comparison with SOTA Methods:
We compare our hybrid approach against two relevant SOTA baselines on the ROAD dataset to further contextualize our method’s performance. Our proposed hybrid method integrates structural graph features with temporal dynamics to enhance anomaly detection. SOTA baselines include:
Moriano et al. [moriano2024benchmarkingunsupervisedonlineids] — An unsupervised time series approach which is evaluated on the ROAD dataset [verma2022addressing] and reports AUC-ROC and TTW using sliding window parameters $(\omega_{t},\delta_{t})$ aligning with our evaluation design. We replicated their best-performing configuration and included results for each masquerade attack.
CANShield [shahriar2022canshield] — A deep learning-based method that operates on signal-level representations, applied to the ROAD dataset [verma2022addressing]. While effective in detecting various attack classes, it does not report TTW nor explicitly define $(\omega_{t},\delta_{t})$ values. Instead, it segments data internally using window_step_test parameters. Additionally, due to the high-dimensional signal-level input and deeper architecture, the approach is considerably more computationally intensive on ROAD. As such, TTW is omitted and only AUC-ROC is reported for this baseline.
Table VIII summarizes the attack-wise AUC-ROC and TTW scores. Our approach achieves the highest AUC-ROC (0.99) across all five masquerade attacks while consistently demonstrating lower TTW per window. The optimal $(\omega_{t},\delta_{t})$ values for our configuration are presented in Table VI. For Moriano et al. [moriano2024benchmarkingunsupervisedonlineids], the values correspond to their best-performing public configuration. We show the effectiveness and practicality of our method for real-time detection, highlighting its robust performance under constrained latency budgets. The code to reproduce these baselines has been added to our GitHub repository [GraphML-CONTROLLER-AREA-NETWORK-Attack-Detection].
TABLE VIII: COMPARATIVE PERFORMANCE OF OUR HYBRID APPROACH VERSUS SOTA METHODS ON THE ROAD DATASET. AUC-ROC AND TTW VALUES CORRESPOND TO EACH METHOD’S BEST-PERFORMING CONFIGURATION PER ATTACK.
Attack Name Metric Ours Moriano CANShield correlated_signal AUC-ROC 0.99 0.96 0.93 TTW (s) 0.44 0.67 N/A max_speedometer AUC-ROC 0.99 0.97 0.94 TTW (s) 1.85 2.01 N/A max_engine_coolant AUC-ROC 0.99 0.94 0.91 TTW (s) 0.92 1.04 N/A reverse_light_off AUC-ROC 0.99 0.95 0.92 TTW (s) 0.73 0.89 N/A reverse_light_on AUC-ROC 0.99 0.93 0.90 TTW (s) 1.10 1.36 N/A Average AUC-ROC 0.99 0.95 0.92 TTW (s) 0.97 1.19 N/A
CANShield does not explicitly define $(\omega_{t},\delta_{t})$ values but adopts an equivalent segmentation strategy using window_step_test parameters; TTW is omitted as its framework does not report latency per window.
V-E Evaluation Based on Sample Windows
We investigate the effect of sample-based windows in our detection framework’s performance, as opposed to time-based windows. Our aim is to understand the impact of varying the number of CAN message samples in each sliding window on masquerade attack detection. In this context, time-based windows refer to windows defined by a fixed duration, whereas sample-based windows are defined by a fixed number of CAN messages. We experiment with different $\omega_{s}$ and $\delta_{s}$ to find the optimal configurations for detecting anomalies.
Due to space constraints, we do not include detailed heatmap results in this version of the paper. However, note that our sample-based results suggest that the integration of time series features with graph embeddings enhances the model’s ability to detect anomalies, even when requiring slightly larger sample windows. The improvement in detection capability comes with a measurable increase in computational requirements, as quantified by the TTW metrics in Section V-C. Further optimization and hardware-specific tuning will be necessary to ensure real-time feasibility in production vehicles.
We also found the distributions of AUC-ROC values to be significantly different between the two settings. The Mann-Whitney $\mathrm{U}$ and $\mathrm{KS}$ tests confirmed this difference, indicating the superior detection performance of the combined setting. Table IX summarizes the results, showing significant differences across all tested masquerade attacks. For instance, correlated_signal_attack yielded a Mann-Whitney U statistic of 1069.5 with a $p$-value of 9.84e-07, and a KS statistic of 0.58 with a $p$-value of 2.61e-06. Similar trends were observed for other attacks, such as max_engine_coolant_attack and max_speedometer_attack, confirming the effectiveness of incorporating time series features with graph embeddings. The results for each attack type consistently indicate that the combined setting outperforms the embeddings-only setting.
TABLE IX: MANN-WHITNEY U AND KOLMOGOROV-SMIRNOV TEST RESULTS FOR SAMPLE-BASED WINDOWS.
Attack Name U Statistic $p$-value KS Statistic $p$-value correlated_signal 1069.5 9.84e-07 0.58 2.61e-06 max_speedometer 1085.0 4.20e-07 0.56 9.30e-06 max_engine_coolant 811.5 0.033 0.33 0.018 reverse_light_off 1040.5 4.84e-06 0.50 9.36e-05 reverse_light_on 1112.0 8.35e-08 0.61 6.75e-07